# Scalable Private Search with Wally

### v2

@@ 4 Wally Protocol
In this section, we describe the complete Wally protocol for a single epoch with $U$ clients.
@@
The server uses Algorithm 1 to encode the database with $N$ embeddings into $K*d$ vectorized plaintexts. Specifically, the algorithm first divides the input embeddings into $K$ clusters using K-means clustering. We pick $K$ such that the maximum cluster size is less than SHE polynomial dimension $n$ with high probability, which ensures that for each cluster all the embeddings can be encoded into $d$ plaintext vectors. The algorithm then iterates over each cluster separately. Recall each cluster consists of $|C[i]|/d$ column-wise $d*d$ matrices. The algorithm then fills $d$ plaintext vectors for each cluster. In a plaintext vector $j$, the algorithm packs $j$’th diagonal of all $|C[i]|/d$ matrices. Once $d$ plaintext vectors are generated, the algorithm iterates over a group of $g=\sqrt{d}$ plaintexts and rotating each group in increasing multiples of $-g$. Note that these rotations are performed on plaintexts so their cost are negligible.
@@
The client uses Algorithm 2 to generate queries and their schedule at the start of an epoch. The client locally stores the set of cluster centroids $c$ that the server generates during initialization. The algorithm first generates independent queries for at most $\Delta$ real query embeddings. To generate a $j$’th real query, the algorithm picks a centroid $\mathsf{id}_{l}$ nearest to $j$’th query embedding. Then, the algorithm copies the embedding $n/d$ times into a plaintext vector $p$ of size $n$. This is done to take advantage of SHE packing. The algorithm then generates a fresh SHE secret key $\mathsf{sk}$ and evaluation key $\mathsf{evk}$, and encrypts $p$ using $\mathsf{sk}$.
Next, the algorithm generates fake queries. Here, the algorithm iterates over each cluster, sampling a number of fake queries from the negative binomial distribution $\mathcal{NB}(r/U,p)$. To generate a single fake query, the algorithm picks a random centroid, generates a fresh SHE secret key $\mathsf{sk}$ and evaluation keys $\mathsf{evk}$, and encrypts $\mathbf{0}$ using $\mathsf{sk}$.
@@ 5 Security
In this section, we prove that if honest clients follow Algorithm 2 to generate queries, then the overall system achieves $(2\epsilon,2\Delta\delta)$-differential privacy in each epoch. Recall from Algorithm 2 that each query consists of the nearest centroid and the embedding. We only focus on proving the privacy of centroids because the IND-CPA security of SHE protects query embeddings. Therefore, we model each query as a nearest centroid without an embedding in this section. During each epoch, the server receives multiple queries for each cluster. Towards the end of epoch, the server’s view can be considered as a noisy histogram over all clusters. We consider this noisy histogram as an output of the DP mechanism.
@@ 5.1 DP Security in the Central Model
Basic composition also yields that the composed protocol has $\delta^{\prime}:=2\Delta\delta$ since at most $2\Delta$ clusters differ.
We emphasize that the proof is over a particular epoch $S$. If clients participate in many epochs, say $l$, then basic composition yields $(2l\epsilon,2l\cdot\Delta\delta)$-DP.
@@ 5.2 Generating view of the server
6 Optimizations
@@
Here we present our optimizations for semantic and keyword search. Our optimizations revolve around the [Bra12, FV12] and BGV [BGV14] SHE schemes. Further, we describe several optimizations to the MulPIR scheme [ALP+21]. First, we describe the standard RNS optimization in SHE [GHS12] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [HLHD22].
@@
We use Cheetah’s [HLHD22] method to compress ciphertexts after homomorphic computation: we modulus switch down to the smallest RNS modulus, $Q^{\prime}=q_{0}$, then drop the least significant bits of the ciphertext. This technique compresses the response size further than simply modulus-switching down, but it also adds noise. Cheetah incorrectly models the ciphertext as noiseless (Appendix F in [HHC+23]), but we derive the proper analysis as follows. Dropping the LSBs of a BFV ciphertext $\mathsf{ct}=(c_{0},c_{1})$ can be analyzed by expressing the polynomials as high and low bit polynomials $c_{0}=2^{b_{0}}c_{0}^{h}+c_{0}^{l},\ c_{1}=2^{b_{1}}c_{1}^{h}+c_{1}^{l}$. Then, the new decryption equation is
@@
6.1 Approximate Nearest Neighbors Optimizations
@@
Wally uses the “baby-step giant-step” (BSGS) optimization [HS18] to minimize the number of rotation keys: for $d$-dimensional inner products, BSGS requires $2\sqrt{d}$ rotations for step sizes of $1$ (baby) and $\sqrt{d}$ (giant). See Algorithm 1 in [HS18] for the BSGS algorithm. Additionally, we are able to use two rotation keys since $d<n/2$ in our use case. This optimization leads us to a query size of one ciphertext $\mathsf{ct}$ and two rotation keys $\mathsf{evk}=(\mathsf{evk}_{1},\mathsf{evk}_{\sqrt{d}})$. For key-switching, and relinearization in PIR below, we use the hybrid-GHS [GHS12] key-switching strategy described in-detail by Kim et al. [KPZ21].
Say we require a plaintext modulus of $t\approx 2^{29}$ for $\sim 14$ bits of plaintext precision. The smallest BFV ring dimension supporting this is $n=8192$ which requires a evaluation key modulus of $Q\approx 2^{165}$, ciphertext modulus of $Q^{\prime}\approx 2^{110}$ for 128 bits of post-quantum security [APS15] [^4] https://github.com/malb/lattice-estimator. A query consists of $2$ polynomials for the ciphertext and $2$ polynomials for the evaluation keys, or $225$kB and $676$kB respectively, for a total query size of $901$kB for these parameters.
@@
Wally achieves minimal bandwidth per client query by a novel use of the Chinese remainder theorem (CRT) on the BFV plaintext space. The CRT on plaintexts allows us to keep a set amount of plaintext bits while re-using one set of evaluation keys across all plaintext spaces since BFV’s evaluation keys are independent of the ciphertext modulus, unlike BGV evaluation keys. Therefore, we encrypt a vector $\mathbf{\widetilde{v}}$ as $\mathbf{\widetilde{v}}\mod t_{i}$ for NTT-friendly primes $\{t_{i}\}$. The computation is correct if the inner products do not wrap around modulo $t=t_{0}t_{1}\cdots t_{l}$, but the individual ciphertexts can have their computation wrap around modulo each $t_{i}$. Therefore, we can drop the ring dimension $n^{\prime}=4096$ and use two ciphertexts, modulo $t_{0}=40961\approx 2^{15}$ and modulo $t_{1}=65537\approx 2^{16}$ for $t_{0}t_{1}\approx 2^{31}$, half the bandwidth from ciphertexts, and use evaluation keys in dimension $n^{\prime}$, or one fourth the bandwidth from evaluation keys. Concretely, this leaves us with $113$kB for ciphertexts and $170$kB for evaluation keys, $283$kB in total, over a 3×3$\times$3 × improvement in query size without sacrificing plaintext precision.
@@ 6.2 PIR Optimizations
In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication. For example, SEAL [CLP17] on an AWS t2.2xlarge AWS instance has ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $3.6ms$ and ciphertext multiplication is $12.1ms$ [MR23].
@@
MulPIR is an optimized query expansion which originated with Angel et al. [ACLS18]. In short, a query index is encoded in a plaintext polynomial then expanded into $k$ ciphertexts by calling $k$ ciphertext rotation operations (Galois automorphisms followed by key-switching). We noticed that some of these expansions can be substituted with linear operations: call $\mathsf{ct}\pm\mathsf{ct}^{\prime}$ where $\mathsf{ct}^{\prime}$ is $\mathsf{ct}$ rotated. In general, we saw a $10-25\%$ improvement in MulPIR’s expansion from our optimization.
@@
We apply the lazy rescaling technique of Kim et al [KPZ21] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [KPZ21, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.
Handling Very Large Entries
@@
This approach results in $n\cdot t$ ciphertext-plaintext multiplications and $d_{2}\cdot t$ ciphertext-ciphertext multiplications, where $t$ is the number of plaintexts used to encode one entry. The number of ciphertext-ciphertext multiplication grows linearly with $t$ and remains a concretely small value when entries are not large. But it will blow up with extremely large entries (e.g. when they are high-resolution photos).
@@
The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext mutliplications to compute, which is independent of the entry size and $t$. This optimization reduce the computation overhead when $t$ is larger than the first dimension $d_{1}$.
@@
BFV’s multiplication step over the integers is done by extending the modulus $Q$ to a larger modulus $PQ$ so the multiplication does not wrap around. We use Bajard et al.’s method to extend the basis, “fast basis conversion” in [BEHZ16] and re-use the basis elements in $Q$ for a $\sim 15\%$ improvement in BGV multiplication.
@@
We add a conditional, lazy modular reduce to Harvey’s NTT [Har14]. The main idea is that we only reduce during the butterfly if the integers would exceeds the machine word size. This saw a $20-25\%$ improvement in forward NTT times and $15-20\%$ improvement in inverse NTT times for $61$-bit primes. Note, we saw no improvement on primes near $64$-bits.
7 Performance
In this section, we demonstrate Wally’s scalability and concrete performance as seen in experiments. We also track Wally’s performance tradeoffs: fake query overhead versus number of clients, fake query overhead versus the DP parameters, and the difference in low-precision performance versus high-precision performance.
Experimental setup.
We ran server-side computations on an AWS c5.24xlarge instance with 192GB of RAM, single-threaded. We extrapolate throughput from the single-threaded times seen on this AWS instance.
End-to-end performance.
Metric Low prec. Mid prec. Tiptoe
Exp. query (MB) 0.67 0.84 16.2
Exp. response (MB) 0.16 0.18 3.2
Queries/s 2998 1499 47
8M queries (hrs) 0.4 1.5 47.5
Fake query 3×3$\times$3 × 3×3$\times$3 × --
overhead
Precision (bits) 6 13 4
Privacy DP DP SHE
Network req. AN AN --
Table 2: Wally performance for a parameter set with BFV dimension $n=4096$ on a 35M entry dataset with the maximum cluster size as $4096$ compared to Tiptoe. Query and response times and sizes for Wally include the number fake query overhead (3×3$\times$3 × or two fake for each real query). Throughput for Wally is measured for 8M real queries, or 24M in total. All Wally responses are computed as the encrypted scores plus $44$kB of metadata.
Table 2 describes our main experiments in comparison with Tiptoe. Here, we use two parameter sets for Wally: first, we use a small BFV parameter set ($n=4096$, $\log_{2}t=15$, and $\log_{2}q=83$) and second, a plaintext CRT (Section 6) parameter set which includes the former. The plaintext precision is $(\log_{2}t)/2-(\log_{2}d)/4$ since the inner-product computation grows heuristically as $\sqrt{d}\|\mathbf{\tilde{e}}\|_{\infty}^{2}$ where $\mathbf{\tilde{e}}=\alpha\cdot\mathbf{e}$ is a scaled (unit) embedding vector, $\|\mathbf{e}\|_{2}=1$. The former parameter set enables six bits of plaintext precision whereas the latter enables 13 bits.
We set $\Delta=2$ messages per client per epoch and calculated the DP overhead by using $(\epsilon,\delta)=(1,2^{-20})$ after the loss in Theorem 2 with four million clients per epoch. The fake query overhead is derived by the expected number of fake queries, for the negative binomial mechanism, $\mathcal{NB}(r,p)$, as
where $K$ is the total number of clusters and $U$ is the total number of clients.
Wally’s ability to cluster the dataset and keep server computation time proportional to the maximum size of a cluster is demonstrated in its small per-query server computation time, $95$ms for each real or fake query, and its relatively small expected query size, $679$kB. We keep each cluster size less than the total number of plaintext slots in each BFV ciphertext, $n=4096$. Otherwise, Wally can compute similarity scores in parallel if the cluster size is larger than $n$. Large meta-data sizes can be retrieved by adding another round of PIR after the similarity computation. The throughput times represent how many queries per second can be processed on $K=8545$ clusters during an epoch needed to serve $24$M total queries, real and fake. This throughput calculation leads to a minimum epoch size of $44$ minutes serving roughly 3k queries per second. This compares to the $47.5$ hours Tiptoe would take to serve $8$M queries.
Comparing to Tiptoe.
Wally has the following two extra assumptions compared to Tiptoe:
  • Wally requires sufficiently many honest clients in the system during an epoch.
  • It also needs an anonymization network that removes all identifying information from incoming messages and randomly permutes messages over a short amount of time, e.g., one second time slots.
In other words, Wally trades off scalability for privacy. Such a tradeoff is needed since oblivious, cryptographically secure systems require the server to compute over the entire database and cannot scale to many clients.
Tiptoe ran its encrypted similarity computation on a cluster of $80$ r5.xlarge AWS instances, each with 4 vCPUs and 32 GB of DRAM, and saw a latency of $11.2$ seconds per query, of which is a significant amount of per-query preprocessing (8.7 seconds). Each server held $8$GB to $12$GB of the plaintext database. We emphasize that each Tiptoe query computes of the entire database (all clusters), and we estimate their total single threaded time as 80×80$\times$80 × their per query time, or $183$ seconds per query. We obtained this estimate by dividing the pre-processing (token) times and the download bandwidths in Table 7 in [HDCZ23] by six since our embedding dimension is 3×3$\times$3 × higher and Tiptoe generates two pre-processing elements per query, “tokens”: one each for a similarity computation and PIR. The non-pre-processing times and bandwidths, as well as the token upload bandwidth, are divided by 3×3$\times$3 × again due to the difference in embedding dimension and that they use the same token upload for both the similarity computation and PIR. Next, the server compute times are multiplied by $80$ since they use $80$ clusters, yielding an estimated $163$ seconds, single-threaded, for a homomorphic similarity computation, i.e., a Wally query. Given this time, we divide it by the number of clusters we expect in Wally, $8545$, to estimate throughput in Table 2.
In comparison, Wally can process each query in parallel since queries, real and fake, have their computation restricted to a single cluster. If there are $U$ clients making at most $\Delta$ queries, then Wally can set a epoch time as the minimal amount of time needed to compute $\phi\times U\times\Delta$ queries. Since each client sends a query at a random time over an epoch, the average response time is half the epoch whereas the worst case time is the epoch length. We derived Wally’s throughput numbers in Table 2 using $95$ms per query, $U=4$M, $\Delta=2$, and $\phi=2$. This leads to $24$M queries in 44 minutes, of which $8$M are real queries, on $K=8545$ clusters.
Scaling DP with $\epsilon$.
Precision.
Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section 6. The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$. Lastly, the 29 precision parameters use two $n=8192$ BFV parameter sets with $\log_{2}t_{1}=29$ and $\log_{2}t_{2}=30$.
Precision (bits) 6 13 29
Query (KB) 226 283 901
Response (KB) 53 63 196
Server comp. (ms) 95 190 580
Scaling with the number of clients.
Number of Clients 0.5M 1M 2M 10M
Fake queries ($\Delta=1$) $7.5$ $4.8$ $2.9$ $1.4$
Exp. MB ($\Delta=1$) $1.9$ $1.08$65.65.65.65 $3.11$
Fake queries ($\Delta=4$) $35.3$ $18$ $9.6$ $2.7$
Exp. MB ($\Delta=4$) $8$ $4.1$ $2.2$61.61.61.61
7.1 Keyword Query Performance
Table 5 shows single-threaded keyword PIR server performance, measured on an M1 Max laptop. Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters. The request size and the response size for each entry in the table is 728KB and 39KB, respectively.
Database Entries Entry Size (bytes) Time (s)
100K 10 0.2
100K 100 0.5
1M 10 0.7
1M 100 2
10M 10 2.7
100M 100 6.3
Table 5: Keyword-PIR, server run times for Wally. These times include the optimizations in Section 6.
8 Conclusion

### v3

@@ 5 Wally Protocol
In this section, we describe the complete Wally protocol for a single epoch in semantic search. We assume that $U$ honest clients are present to query within that epoch. Due to the space limitation, we only describe the semantic search protocol in detail, but keyword search follows a similar protocol with changes mentioned in Section B.
@@
The server uses Algorithm 1 to encode the database with $N$ embeddings into $K*d$ vectorized plaintexts. Specifically, the algorithm divides the input embeddings into $K$ clusters using K-means clustering. The algorithm then iterates over each cluster separately. Each cluster consists of $|C[i]|/d$ column-wise $d*d$ slices. The algorithm then fills $d$ plaintext vectors for each cluster. In a plaintext vector $j$, the algorithm packs $j$’th diagonal of all $|C[i]|/d$ matrices. Once $d$ plaintext vectors are generated, the algorithm iterates over a group of $g=\sqrt{d}$ plaintexts, rotating each group in increasing multiples of $-g$. These rotations are required for the baby-step giant step (BSGS) optimization [27]. Note that these rotations are performed on plaintexts, so their cost are negligible.
@@
Each client uses Algorithm 2 to generate queries and their schedule at the start of an epoch. The client locally has a set of cluster centroids $c$ that the server generates during initialization. The algorithm first generates independent queries for at most $\Delta$ real query embeddings. To generate a $j$’th real query, the algorithm picks a centroid $\mathsf{id}_{l}$ nearest to $j$’th query embedding. Then, the algorithm copies the embedding $n/d$ times into a plaintext vector $p$ of size $n$. This is done to take advantage of SHE packing. The algorithm then generates a fresh SHE secret key $\mathsf{sk}$ and rotation evaluation key $\mathsf{evk}$, and encrypts $p$ using $\mathsf{sk}$.
Next, the algorithm generates fake queries. Here, the algorithm iterates over each cluster, sampling a number of fake queries from the negative binomial distribution $\mathcal{NB}(r/U,p)$. To generate a single fake query, the algorithm picks a random centroid, generates a fresh SHE secret key $\mathsf{sk}$ and evaluation keys $\mathsf{evk}$, and encrypts $\mathbf{0}$ using $\mathsf{sk}$. Note that the client must generate a fresh evaluation key to keep each query anonymous. Also, each evaluation key must include a rotation key for steps $1$ and $g$, because the server computation involves rotating by these steps.
@@ 6 Security
In this section, we prove that if honest clients follow Algorithm 2 to generate queries, then the overall system achieves $(2\epsilon,2\Delta\delta)$-differential privacy in each epoch. Recall from Algorithm 2 that each query consists of the nearest centroid and the embedding. We only focus on proving the privacy of centroids because the IND-CPA security of SHE protects query embeddings. Therefore, we model each query as a nearest centroid without an embedding in this section. During each epoch, the server receives multiple queries for each cluster. Towards the end of the epoch, the server’s view can be considered as a noisy histogram over all clusters. We consider this noisy histogram as an output of the DP mechanism.
@@ 6.1 DP Security in the Central Model
Basic composition also yields that the composed protocol has $\delta^{\prime}:=2\Delta\delta$ since at most $2\Delta$ clusters differ. ∎
We emphasize that the theorem is stated for a particular epoch $S$. If clients participate in many epochs, say $l$, then basic composition yields $(2l\epsilon,2l\cdot\Delta\delta)$-DP.
@@ 6.2 Generating view of the server
7 Implementation and Performance Evaluation
In this section, we demonstrate Wally’s concrete performance via a set of experiments. We also showcase Wally’s performance tradeoffs: fake query overhead versus number of clients, fake query overhead versus the DP parameters, and the difference in low-precision performance versus high-precision performance.
7.1 Implementation
We have implemented Wally in Swift atop the Apple swift-homomorphic-encryption library [7]. Following the library, we use BFV [21] to implement the SHE parts Wally. However, our protocol could easily be implemented using other SHE schemes such as BGV [11].
BFV parameters
The choice of BFV parameters provides a trade-off between computation and communication in Wally. At a high level, larger parameters provide better computation but result in high request size. We tried multiple sets of parameters and picked the one that provided the best performance. We set polynomial modulus $n=4096$, ciphertext modulus $\log_{2}{Q}=83$, and plaintext modulus $\log_{2}{t}=15$ bits. These parameters allow a precision of six bits. To increase the precision, we use plaintext CRT optimization described in Section B. Recall that this optimization’s key benefit is increasing precision without increasing the evaluation key size (which is a dominating factor in each query). For this we set first plaintext modulus $\log_{2}{t_{0}}=15$ bits and second plaintext $\log_{2}{t_{1}}=16$. This gives us a precision of 13 bits.
Differetnial privacy parameters
Differential privacy parameters provide a trade-off between privacy and performance overhead due to fake queries. For a fixed failure probability $\delta$, smaller $\epsilon$ provides stronger privacy but worse performance because each client is required to make more fake queries. We set $\delta=2^{-20}$, which means DP will fail only once a million and $\epsilon=1$, providing a decent privacy guarantee.
7.2 Performance Evaluation
Experimental setup.
We ran server-side computations on an Intel Xeon w3-2423 instance with 32GB of RAM and 6 cores.
End-to-end performance.
Metric Low prec. Mid prec. Tiptoe
Query Size (MB) 0.67 0.84 16.2
Response Size
(MB) 0.16 0.18 3.2
Time to process
queries 39 mins 44 mins 2.8 years
Precision (bits) 6 13 4
Privacy DP DP SHE
Table 2: Performance comparison of Wally and Tiptoe for database of 35 millions entries. Query and response times and sizes for Wally include the number fake query overhead (3×3$\times$3 × or two fake for each real query). Throughput for Wally is measured for 8 million real queries, or 24 million in total. All Wally responses are computed as the encrypted scores plus $44$ KB of metadata.
Table 2 describes our main experiments compared to Tiptoe. The server database consists of 35 million entries divided into around 8500 clusters. We keep each cluster size less than the number of plaintext slots in each BFV ciphertext, $n=4096$. We assume each client makes $\Delta=2$ queries per epoch for these experiments. Moreover, there are four million honest clients present in an epoch. This results in eight million real queries in an epoch.
For Wally, we calculated the overhead by using Theorem 2. The fake query overhead is derived by the expected number of fake queries for the negative binomial mechanism, $\mathcal{NB}(r,p)$, as
where $K$ is the total number of clusters and $U$ is the total number of clients. For our selected parameters, for each real query, the client is required to make just under two fake queries. Therefore, the total number of queries within an epoch is 24 million. In Table 2, we included fake query overhead in query request, response and total computation in Wally . Epoch size in Wally equals the total time it takes to process all the queries (real and fake).
For Tiptoe, we assume there are eight million queries the server processes one by one. Tiptoe requires running experiments on real compute nodes; due to resource limitations, we carefully estimate their results using Table 7 in [29].
Before discussing Wally’s performance, we emphasize that Wally requires the following two extra assumptions not required in Tiptoe.
  • Wally requires sufficiently many honest clients during an epoch.
  • It also needs an anonymization network that removes all identifying information from incoming messages and randomly permutes messages over a short time, e.g., one-second time slots.
In other words, Wally trades off scalability for privacy.
Wally has significantly better performance than Tiptoe across all the metrics. Specifically, the query size (including overhead due to fake queries) in Wally 19-23 times smaller than in Tiptoe. This is because the query in Wally only covers a single cluster, while Tiptoe covers all clusters. Similarly, the response size is 3.1-3.6 smaller. Wallyś efficient server computation makes it an ideal candidate for practical deployments in comparison to Tiptoe. For example, to process eight million queries, the server with 320 cores requires around 39 minutes in Wally . In Tiptoe, the same server would require more than two years to process these queries, which is unsuitable for most practical search services.
Scaling DP with $\epsilon$.
Precision.
Table 3 shows how each query grows as we increase precision, i.e., as we increase the BFV parameters using our plaintext CRT optimization described in Section B. The six bit precision parameters are the same as we used in Table 2 while the $13$ bit parameters use the plaintext CRT on $\log_{2}(t_{1})=15.3$ and $\log_{2}(t_{2})=16$, $\log_{2}(t)=\log_{2}(t_{1}t_{2})=31.3$. Lastly, the 29 precision parameters use two $n=8192$ BFV parameter sets with $\log_{2}t_{1}=29$ and $\log_{2}t_{2}=30$.
Precision (bits) 6 13 29
Query (KB) 226 283 901
Response (KB) 53 63 196
Server comp. (ms) 19 35 110
Scaling with the number of clients.
Number of Clients .5M 1M 2M 10M
Fake queries ($\Delta=1$) $7.5$ $4.8$ $2.9$ $1.4$
Exp. MB ($\Delta=1$) $1.9$ $1.08$ $0.65$ $3.11$
Fake queries ($\Delta=4$) $35.3$ $18$ $9.6$ $2.7$
Exp. MB ($\Delta=4$) $8$ $4.1$ $2.2$ $0.61$
7.3 Keyword Query Performance
Table 5 shows single-threaded keyword PIR server performance, measured on an Intel Xeon w3-2423 CPU with 32GB of RAM using swift-homomorphic-encryption [7]. Each entry used a BFV parameter set of $n=4096$, $\log_{2}Q=83$ with three ciphertext moduli, and $\log_{2}t=6$ using MulPIR [4]. Each response size was 27KB and the query sizes were 368 KB, 539 KB, 539 KB, and 624 KB for each respective row in the table. Note, however, that each dataset size requires manual tuning for optimal encryption and cuckoo hashing parameters.
Database entries Entry Size (Bytes) Time (ms)
100k 10 166
100k 100 465
1 million 10 565
1 million 100 1700
Table 5: Keyword-PIR, or exact query, server run times for Wally. These times include the optimizations in Section B.
8 Conclusion
9 Open science and ethics discussion
9.1 Open Science Policy
We have ensured open science policies by taking the following steps: We develop our system using an open-source library. The paper discusses and provides all the research techniques we used to develop the end system. We provide open access to the paper by sharing it on Arxiv Once accepted, we will reveal part of the code that will allow reproducibility. The entire protocol is a simple extension of the code we will open source, but we cannot provide a full implementation of the end-to-end deployed system due to our company’s IP policy.
9.2 Ethics
This paper complies with all the ethics guidelines. Specifically, we do not use private data or code without developers’ permission or run experiments with people. Our protocol protects user privacy in search systems without affecting the system’s usability; therefore, it benefits all parties. We also took essential disclosures from our employer to share the results given in our paper.
Appendix A Related Work
Protocols like Tor [16] and iCloud Private Relay [6] do not provide query privacy. Further, the server can learn the distribution of queries to each cluster when these protocols are applied to clustered databases.
Recently, there has been a number of works making private search engines using homomorphic encryption, and potentially PIR, in order to obtain full cryptographic obliviousness. Tiptoe [29], the most recent and performant, has been discussed above in-detail. Other, less efficient solutions, include Coeus [1] as well as Hers [20], with the latter providing privacy for the server’s database as well. Adjusting Wally to provide security for the server would decrease performance since we would need to add circuit privacy to the server’s SHE computation, raising the RLWE parameters to the next largest ring dimension $n$ for noise flooding[8, 35], and including a client public key in each query [23].
Another related area of work is differentially-private multi-server PIR schemes [38, 2]. In both works, servers learn differentially private, noised histograms over queries. The former, [38], has clients generate fake queries whereas the latter, [2], has the servers inject fake queries. Both require a secure multiparty computation over the entire database. Notably, [2] can serve one million queries in about two seconds on a 100K entry database with eight servers in a dishonest majority security model and a ten minute pre-computation. Their offline and online computation grows linearly with the size of the server’s database.
Differential privacy has also been applied to multiple anonymous messaging systems [41, 39, 34]. These are large (many server) mix-net-like protocols where the servers add in fake queries to hide user messages with a differential privacy guarantee. Though they are for a different application, their use of differential privacy is similar to Wally’s.
Appendix B Optimizations
@@
Here we present our optimizations for both use cases: approximate nearest neighbor search and exact retrieval (PIR). Our optimizations revolve around the [10, 21] and BGV [11] SHE schemes. Further, we describe several optimizations to the MulPIR scheme [4]. First, we describe the standard RNS optimization in SHE [24] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [31].
@@
We use Cheetah’s [31] method to compress ciphertexts after homomorphic computation: we modulus switch down to the smallest RNS modulus, $Q^{\prime}=q_{0}$, then drop the least significant bits of the ciphertext. This technique compresses the response size further than simply modulus-switching down, but it also adds noise. Cheetah incorrectly models the ciphertext as noiseless (Appendix F in [30]), but we derive the proper analysis as follows. Dropping the LSBs of a BFV ciphertext $\mathsf{ct}=(c_{0},c_{1})$ can be analyzed by expressing the polynomials as high and low bit polynomials $c_{0}=2^{b_{0}}c_{0}^{h}+c_{0}^{l},\ c_{1}=2^{b_{1}}c_{1}^{h}+c_{1}^{l}$. Then, the new decryption equation is
@@
B.1 Approximate Nearest Neighbors Optimizations
@@
Wally uses the “baby-step giant-step” (BSGS) optimization [27] to minimize the number of rotation keys: for $d$-dimensional inner products, BSGS requires $2\sqrt{d}$ rotations for step sizes of $1$ (baby) and $\sqrt{d}$ (giant). See Algorithm 1 in [27] for the BSGS algorithm. Additionally, we are able to use two rotation keys since $d<n/2$ in our use case. This optimization leads us to a query size of one ciphertext $\mathsf{ct}$ and two rotation keys $\mathsf{evk}=(\mathsf{evk}_{1},\mathsf{evk}_{\sqrt{d}})$. For key-switching, and relinearization in PIR below, we use the hybrid-GHS [24] key-switching strategy described in-detail by Kim et al. [32].
Say we require a plaintext modulus of $t\approx 2^{29}$ for $\sim 14$ bits of plaintext precision. The smallest BFV ring dimension supporting this is $n=8192$ which requires a evaluation key modulus of $Q\approx 2^{165}$, ciphertext modulus of $Q^{\prime}\approx 2^{110}$ for 128 bits of post-quantum security [3] [^3] https://github.com/malb/lattice-estimator. A query consists of $2$ polynomials for the ciphertext and $2$ polynomials for the evaluation keys, or $225$kB and $676$kB respectively, for a total query size of $901$kB for these parameters.
@@
Wally achieves minimal query bandwidth per client query by a novel use of the Chinese remainder theorem (CRT) on the BFV plaintext space. The CRT on plaintexts allows increase the plaintext precision without increasing evaluation key size since BFV’s evaluation keys are independent of the plaintext modulus. Therefore, we encrypt a vector $\mathbf{\widetilde{v}}$ as $\mathbf{\widetilde{v}}\mod t_{0}$ and $\mathbf{\widetilde{v}}\mod t_{1}$ for NTT-friendly primes $\{t_{0},t_{1}\}$. The computation is correct if the inner products do not wrap around modulo $t=t_{0}t_{1}$, but the individual ciphertexts can have their computation wrap around modulo each $t_{i}$.
@@ B.2 PIR Optimizations
In general, the cost of a ciphertext-ciphertext multiplication is much higher than a ciphertext rotation and both are much higher than addition or plaintext multiplication. For example, SEAL [13] on an AWS t2.2xlarge AWS instance has ciphertext addition and plaintext multiplication as less than $100\mu s$ while ciphertext rotation is $3.6ms$ and ciphertext multiplication is $12.1ms$ [37].
@@
MulPIR is an optimized query expansion which originated with Angel et al. [5]. In short, a query index is encoded in a plaintext polynomial then expanded into $k$ ciphertexts by calling $k$ ciphertext rotation operations (Galois automorphisms followed by key-switching). We noticed that some of these expansions can be substituted with linear operations: call $\mathsf{ct}\pm\mathsf{ct}^{\prime}$ where $\mathsf{ct}^{\prime}$ is $\mathsf{ct}$ rotated. In general, we saw a $10-25\%$ improvement in MulPIR’s expansion from our optimization.
@@
We apply the lazy rescaling technique of Kim et al [32] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [32, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.
Optimization for Very Large Entries
@@
This approach results in $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications and $d_{2}\cdot t$ ciphertext-ciphertext multiplications, where $t$ is the number of plaintexts used to encode one entry. The number of ciphertext-ciphertext multiplication grows linearly with $t$ and remains a concretely small value when entries are not large. But it will blow up with extremely large entries (e.g., when they are high-resolution photos).
@@
The outer-product takes $d_{1}\cdot d_{2}$ ciphertext-ciphertext mutliplications to compute, which is independent of the entry size. While the inner products still takes $d_{1}\cdot d_{2}\cdot n$ ciphertext-plaintext multiplications. Therefore, this optimization can reduce the computation overhead when the entries size $t$ is larger than the first dimension size $d_{1}$.
@@
BFV’s multiplication step over the integers is done by extending the modulus $Q$ to a larger modulus $PQ$ so the multiplication does not wrap around. We use Bajard et al.’s method to extend the basis, “fast basis conversion” in [9] and re-use the basis elements in $Q$ for a $\sim 15\%$ improvement in BGV multiplication.
@@
We add a conditional, lazy modular reduce to Harvey’s NTT [28]. The main idea is that we only reduce during the butterfly if the integers would exceeds the machine word size. This saw a $20-25\%$ improvement in forward NTT times and $15-20\%$ improvement in inverse NTT times for $61$-bit primes. Note, we saw no improvement on primes near $64$-bits.